# Unraveling the Geometry of Visual Relational Reasoning

### v1

@@ Introduction
We propose that the trade-off between the generalization signal and D arises from a structured compression effect. All class manifolds—including those for held-out rules—gradually align with a low-dimensional subspace defined by the principal components of the training rules. Across models, the representations of the 35 training rules consistently occupy a low-dimensional space. Notably, we observe a drop in variance explained by the top principal components (PCs) at exactly $35$, matching the number of training rules (figure S4). We hypothesize that the network amplifies the generalization signal by aligning all representations, including those of held-out classes, within this low-dimensional space.
To test this hypothesis, we measured the fraction of variance explained by the top $35$ principal components (PCs) of the training rules for each held-out rule manifold. Higher values indicate more substantial alignment with the low-dimensional subspace of training rules. We observed that as the network signal increases, this alignment becomes more pronounced across models (Fig. 6E-H).
@@
Here, $SNR_{a}$ represents the estimated SNR ( Equation 1) for the ordered rule pair $(a,b)$, averaged across all possible rule pairs in the batch.
To assess the effectiveness of the proposed SNR loss, we trained the top-performing SCL model using this objective and evaluated its generalization across the same $15$ rule splits previously used for training with cross-entropy loss. The model effectively optimized the training rule representations, achieving an average SNR of $9.66\pm 0.22$ (mean±$\pm$±s.e.m.) across training rule pairs, as evaluated on novel test rows of the training rules. More importantly, the model effectively separated held-out rules, yielding an averaged generalization error of $0.162\pm 0.018$ (mean±$\pm$±s.e.m.) (Fig. 7A). This performance was statistically indistinguishable from models trained with cross-entropy (CE) loss using one-hot rule labels (Wilcoxon signed-rank test: statistic = $48.0$, $p=0.26$, n.s.), confirming that SNR loss achieves competitive generalization while structuring learned representations according to theoretical predictions.
To further evaluate the efficacy of SNR loss, we compared its performance against Prototypical loss, a widely used loss function designed to learn structured representations that generalize to unseen classes  (?). Prototypical loss operates by embedding inputs into a high-dimensional representation space and forming class prototypes, defined as the centroids of the representations corresponding to each class. The loss function then encourages representations of samples from the same class to cluster around their respective prototypes. Additional details on Prototypical loss are provided in the Supplementary Text. We trained the SCL model using Prototypical loss on the same $15$ rule splits. The resulting model exhibited a generalization error of $0.227\pm 0.010$, significantly higher than that obtained with SNR loss (Fig. 7A). These findings suggest that while both loss functions promote structured representations, SNR loss is more effective in shaping a geometric representation that enhances generalization.
@@ Discussion
We address the fundamental question of how neural networks discover unseen rules when learning abstract relational rules. We hypothesize that networks trained to perform reasoning tasks generate good representations even for unseen rules. The utility of these representations can be quantified by the geometry of the rule manifolds defined by all representations associated with a given rule. Extending the geometric theory of object manifold  (?) to abstract rules, we predicted an explicit relation between the generalization performance in visual reasoning and key geometric properties: manifold radii, dimensionality, the distance between manifold centroids, and the overlap between these distances and their subspaces of variation.
To test whether the geometric theory developed for object recognition can be applied to visual reasoning tasks, we introduce SimplifiedRPM, a subset of Raven’s Progressive Matrices (RPM), whose simpler design allows a systematic study of generalization ability in visual reasoning and direct comparison with the above theory. Using this dataset, we trained state-of-the-art deep network architectures (ResNet  (?), WReN  (?), ViT  (?), SCL  (?)). We demonstrated the superiority of the SCL (Scatter Compositional Learner) architecture, consistent with previous reports  (?). We then conducted an in-depth comparison between the performance of the SCL network and human participants on the same task. Surprisingly, we found excellent agreement between human and network performance, both in terms of overall performance but also across different abstract rules.
Leveraging on a geometrical theory of manifolds initially applied to object recognition tasks  (?), we show that interpretable representation geometries can explain models’ performance. Further analysis reveals how different architectures balance signal and dimensionality to achieve relational reasoning while identifying bottlenecks limiting their performance. Building on these insights, we propose SNRloss, a theoretically grounded objective function that explicitly optimizes representation geometry. This novel loss function provides adequate training and generalization capabilities. Our work establishes a geometrically informed framework for analyzing neural networks’ capacity to recognize and generalize abstract relations.
A key insight from our study is that neural networks do not generalize equally well across all abstract relations. While prior research has examined abstract relation generalization in neural networks  (?), our work offers a more fine-grained empirical evaluation. Notably, the top-performing model, SCL, aligns closely with human difficulty rankings, with logical rules being the hardest. In contrast, the other three models struggle to generalize arithmetic rules despite achieving high test accuracy. Furthermore, we find that signal, rather than dimensionality or signal-noise overlap, is the primary factor governing these generalization differences among abstract relations. These insights suggest that future models could benefit from strengthening signals for specific relational types.
The best-performing model, SCL, shows a sharp improvement in relational separability only in the last two MLP layers of the network. The earlier layers exhibit only minor gains despite the network being trained end-to-end. This observation suggests that high-level relational abstractions do not gradually build up over the model’s layers but instead emerge abruptly during the final transformations. Geometrically, the signal rises even later in the last MLP layer, while the dimensionality shows a non-monotonic trend—first expanding and then decreasing. The same geometric trend was previously observed in vision models, including ResNet-50, VGG19, and ResNet-152, trained on the ImageNet object classification task  (?). The consistency of these geometric trends across models trained on relational reasoning and object classification tasks suggests that they are not merely artifacts of specific architectures or tasks but may reflect a general mechanism of hierarchical abstraction. In contrast, ViTs show a more gradual refinement of relational separation and a steady signal increase across network layers. This indicates that attention mechanisms may support relational learning differently than convolution-based architectures such as SCL and ResNet.
@@
According to the geometrical theory, higher dimensionality should be favorable for generalization. On the other hand, prior studies suggest that low-dimensional manifolds offer advantages for object recognition, enabling the classification of a greater number of objects  (?, ?). This contrast implies that dimensionality should not be universally maximized or minimized but instead tuned to task demands. Our findings extend this idea by showing that generalization depends not only on dimensionality but mostly on its interaction with pairwise manifold signals. Specifically, we demonstrate that deep networks systematically compress representations into structured, low-dimensional spaces—even for unseen rules. Moreover, this phenomenon is consistently observed across diverse architectures (e.g., ResNet, SCL), indicating that it is a fundamental learning property rather than an artifact of any specific architecture. We observe that late layers play a crucial role in compressing relational representations, a phenomenon consistent with prior work showing that deep networks tend to refine and condense information in later stages of learning  (?). These findings open new avenues to explicitly leverage this signal versus dimensionality trade-off to improve generalization.
@@
Among the tested models, SCL best replicates the human-like difficulty ordering across rules. While prior research has relied on explicit relational priors to achieve human alignment  (?), our results demonstrate that SCL, when trained end-to-end on the task, exhibits human-like performance without requiring such priors. Furthermore, SCL’s generalization errors closely correlate with human response times. This raises a key question: What insights does SCL offer for understanding the brain’s relational reasoning mechanisms? The SCL network parallels the brain’s hierarchy: its CNN-based feature extractor mimics the ventral visual cortex, capturing basic features like shape and color  (?, ?, ?, ?), while its later reasoning module reflects higher cortical areas supporting relational reasoning  (?, ?, ?, ?). We propose that the brain encodes rules of input rows through distinct, rule-specific neural manifolds. Our observation of layer-wise changes in the geometry of these rule manifolds offers hypotheses into how neural representations evolve across brain regions  (?, ?). According to the SCL model, early visual areas exhibit high representational dimensionality (D) but low signal, whereas later regions show larger signals with reduced dimensionality. These predictions can be empirically tested by measuring neural activity during abstract reasoning tasks  (?, ?, ?, ?).
@@
This study aimed to study the generalization of abstract relations. While the humans in our experiment were not explicitly trained on the rules, their reasoning abilities could be considered ”pre-trained” through life experiences, as they may have previously encountered similar rules. In contrast, neural networks were trained under strictly controlled rule splits, ensuring they had no prior exposure to the rules evaluated for generalization. To address this discrepancy, a crucial next step is to develop a new benchmark for human reasoning that incorporates abstract relations unlikely to have been encountered. Such a benchmark would disentangle relational generalization from prior experience, providing a clearer view of their distinct roles in human abstract reasoning. This approach enables a direct comparison between humans and neural networks, offering deeper insights into their generalization mechanisms.
@@
Figure 1: The SimplifiedRPM dataset helps evaluate neural networks in relational reasoning and generalization. (A) Example SimplifiedRPM trial with “constant shape“ rule. Choice 1 is correct; Choice 2 follows “constant number” and is incorrect. (B) We evaluate four representative models on the SimplifiedRPM task: ResNet-50 (orange)  (?), WReN (green)  (?), ViT (purple)  (?), and SCL (yellow-green)  (?). Each model encodes relational rules in rows using high-dimensional relational representations (red) and is trained end-to-end. A linear classifier is appended during training to categorize relational rules. See Methods for details. (C) Test errors for four models across 15 random splits of training and held-out rules, with error bars denoting s.e.m. All models perform significantly above chance ($0.97$) based on a one-sample t-test: ResNet ($t(14)=-312.77$, $p=1.2\times 10^{-28}$), WReN ($t(14)=-343.78$, $p=3.4\times 10^{-29}$), ViT ($t(14)=-388.11$, $p=6.2\times 10^{-30}$) and SCL ($t(14)=-542.53$, $p=5.7\times 10^{-32}$). Pairwise Wilcoxon tests confirm SCL achieves the lowest test error, followed by ViT, WReN, and ResNet ($p$-values: SCL vs. ViT, $3.0\times 10^{-5}$; ViT vs. WReN, $1.6\times 10^{-3}$; WReN vs. ResNet, $3.0\times 10^{-5}$). (D) Schematic illustrating distances (red arrows) between sample and choice row relational representations. The trial is solved correctly if the correct choice row representation is closer to the sample row than the incorrect choice. (E) Generalization error of four models on SimplifiedRPM, averaged over $15$ random training and held-out rule splits. Error bars denote s.e.m. All models perform significantly above chance ($0.5$) based on a one-sample t-test: ResNet ($t(14)=-8.93$, $p=1.8\times 10^{-7}$), WReN ($t(14)=-26.39$, $p=1.2\times 10^{-13}$), ViT ($t(14)=-33.45$, $p=4.6\times 10^{-15}$), and SCL ($t(14)=-19.35$, $p=8.3\times 10^{-12}$). Pairwise Wilcoxon tests confirm SCL achieves the lowest error, followed by ViT, while WReN and ResNet show similarly high errors ($p$-values: SCL vs. ViT, $6.1\times 10^{-5}$; ViT vs. WReN, $3.0\times 10^{-5}$). Asterisks indicate statistical significance: ** ($p<0.01$), *** ($p<0.001$).
Figure 2: SCL Captures Human-like Differential Relational Rule Difficulties. (A-D) Generalization errors of four models on the SimplifiedRPM task, evaluated across three abstract relation types: Progression (blue), Arithmetic (green), and Logical (red). The chance level is 0.5. Each bar represents the average error for trials where the correct rule of the sample row fell into one of these relation types, with error bars denoting s.e.m. (A) ResNet exhibited significantly higher errors for arithmetic rules than logical rules (p = 0.019, Pairwise Wilcoxon test). (B) WReN showed significantly higher errors for arithmetic rules than both progression ($p=4.2\times 10^{-6}$) and logical ($p=9.1\times 10^{-5}$) rules. (C) ViT exhibited higher errors for arithmetic rules compared to both progression ($p=3.8\times 10^{-17}$) and logical ($p=7.7\times 10^{-7}$) rules. Additionally, logical rules had significantly higher errors than progression ($p=2.0\times 10^{-12}$). (D) SCL progression was lower than arithmetic ($p=2.1\times 10^{-17}$) and logical ($p=3.2\times 10^{-8}$). Statistical significance was assessed using Pairwise Wilcoxon tests. Asterisks indicate statistical significance: * ($p<0.01$), *** ($p<0.001$). (E) Comparison of the SCL model’s generalization errors against human performance on the SimplifiedRPM task across 13 selected rule pairs (chance level = 0.5). Each dot represents the average model (x-axis) or human (y-axis) error over 10 questions for a given rule pair, with error bars indicating the s.e.m. The dashed line represents the line of identity. (F) Correlation between the SCL model’s generalization error (x-axis) and human response times (y-axis) for the 13 selected rule pairs. Each dot represents the average model error and corresponding human response time for a given rule pair, with error bars indicating the s.e.m.
Figure 3: Geometric Theory of Rule Manifolds Accurately Predicts Error in the SimplifiedRPM Tasks. (A) Illustration of rule manifolds in the relational representation space. Orange axes denote representation dimensions. Each rule manifold consists of relational representations for rows that follow a specific rule. Two rule manifolds are shown: blue (rule a) and green (rule b) ellipsoids. The two black dots in rule a manifold represent example rows following rule A (e.g., constant shape). Each manifold is characterized by a centroid ($\mathbf{x}_{0}^{a}$ for rule a, blue star; $\mathbf{x}_{0}^{b}$ for rule b, green star), representing the mean position of all relational representations following that rule. Variation is characterized by principal axes ($\mathbf{u}_{i}^{a}$ for the blue rule manifold a and $\mathbf{u}_{i}^{b}$ for the green manifold b), shown as blue and green dashed arrows. Here, $i=1,..,N$, where $N$ is the relational representation dimension. The extent of variation along each axis is quantified by the corresponding radii $R_{i}^{a}$ and $R_{i}^{b}$. Purple line represents signal $\Delta\mathbf{x}_{0}$. (B-E) Comparison of theoretical predictions (red lines) and empirical generalization errors (dots) across four models. Each dot represents a held-out rule pair, where the theoretical SNR predictions are computed using geometric measures (Equation 1), and the empirical generalization errors are averaged over 500 SimplifiedRPM trials. Each plot includes rule pairs from 15 distinct held-out rule splits. Each split consists of five held-out rules, resulting in 20 rule pairs per split. (F-I) Generalization SNR, signal, D, and signal-noise overlap (normalized by signal magnitude) for model relational representations on the held-out rules. The gray bars represent the untrained models’ representation at initialization, while the colored bars correspond to different models: orange for ResNet, green for WReN, purple for ViT, and yellow-green for SCL. Each bar shows the mean value averaged across 15 distinct held-out rule splits, with error bars indicating the s.e.m. Pairwise Wilcoxon tests assess whether the trained model terms differ from the untrained models. Asterisks indicate statistical significance: * ($p<0.01$), *** ($p<0.001$).
Figure 4: Generalization error differences across abstract relation types were primarily influenced by signal rather than dimensionality. (A-D) Generalization signal and (E-H) dimensionality for different abstract relation types—progression, arithmetic, and logical—across four models. Results are averaged over 15 random splits of training and held-out rules. Each bar represents the average error for trials where the correct rule of the sample row belonged to each relation type, with error bars indicating the s.e.m. Statistical comparisons were conducted using the pairwise Wilcoxon test.
Figure 5: Layer-wise Geometric Analysis Reveals Model-Specific Generalization Strategies. (A-D) Generalization error across layers for each model. The plotted values represent 15 distinct held-out rule splits, each containing five held-out rules, leading to 20 rule pairs per split. Error bars denote the s.e.m. across the 15 splits. Representations at each layer are randomly projected to a fixed dimensionality of $N=400$ before computing errors and performing geometric analysis. Generalization errors for each rule pair are evaluated using 500 SimplifiedRPM trials. (E-H) Generalization signal-to-noise ratio (SNR) across layers. (I-L) Generalization signals across layers. (M-P) Generalization dimensionality (D) across layers.
Figure 6: Structured Compression Drives the Trade-off Between Generalization Signal and Dimensionality. (A-D) Generalization signal vs. dimensionality (D) across training epochs for all models. Each point shows average signal and D over 15 held-out splits (five rules per split). Dot size and color indicate training epoch; later epochs are larger, brighter. Spearman’s rank correlation between the generalization signal and D reveals a strong negative correlation across the 15 rule splits for each model (ResNet: $\rho=-0.12\pm 0.12$; WReN: $\rho=-0.41\pm 0.11$; ViT: $\rho=-0.37\pm 0.19$; SCL: $\rho=-0.77\pm 0.05$; mean $\pm$ s.e.m.). This effect was statistically significant, as confirmed by Stouffer’s test (ResNet: $Z=3.41$, $p=3.2\times 10^{-4}$; WReN: $Z=2.38$, $p=8.5\times 10^{-3}$; ViT: $Z=14.76$, $p=1.1\times 10^{-49}$; SCL: $Z=13.04$, $p=3.5\times 10^{-39}$). (E-H) Alignment of held-out rule manifolds with the low-dimensional subspace of training rules across training epochs for all four models. Each point shows the variance explained by the top 35 principal components of training rule representations for a held-out rule manifold, averaged over 15 splits. The size and color of the dots indicate the training epoch, with later epochs depicted by larger, brighter dots. Spearman’s rank correlation between the generalization signal and subspace alignment shows a strong positive correlation across the 15 rule splits for each model (ResNet: $\rho=0.01\pm 0.13$; WReN: $\rho=0.29\pm 0.11$; ViT: $\rho=0.38\pm 0.19$; SCL: $\rho=0.90\pm 0.04$; mean $\pm$ s.e.m.). This effect was statistically significant in all four models except WReN, as confirmed by Stouffer’s test (ResNet: $Z=3.35$, $p=3.9\times 10^{-4}$; WReN: $Z=0.37$, $p=0.35$; ViT: $Z=12.09$, $p=6.0\times 10^{-34}$; SCL: $Z=19.54$, $p=2.5\times 10^{-85}$).We begin plotting from the training epoch, where the structured low-dimensional representation of the training rule first emerges, indicated by the stabilization of its relational representation’s dimensionality. The full trajectory of the generalization signal, D, and alignment is presented in figure S5.
Figure 7: Theory-Grounded SNR Loss Yields Geometry-Balanced Representations. (A) Generalization error for models trained with SNR loss, cross-entropy (CE) loss, and Prototypical (Proto) loss, averaged over 15 rule splits. Error bars denote s.e.m. Pairwise Wilcoxon tests confirm Prototypical achieves the highest error ($p$-values: SNR Loss vs. Proto Loss, $5.1\times 10^{-3}$; CE vs. Proto Loss, $1.0\times 10^{-3}$). (B,C) Comparison of theoretical predictions (red lines) and empirical generalization errors (dots) for the SNR and Prototypical Loss. Each dot represents a held-out rule pair. (D-F) Generalization signal, dimensionality (D), and signal-noise overlap (normalized by signal magnitude) for models trained with SNR, CE, and Proto Loss. Each bar represents the mean value across 15 rule splits, with error bars indicating the s.e.m. Pairwise Wilcoxon tests assess the statistical significance of differences between loss functions. Asterisks indicate statistical significance: * ($p<0.05$), ** ($p<0.01$), *** ($p<0.001$).
Acknowledgments
Funding:
National Institutes of Health grant R01EY026025 (GK)
Swartz Foundation (HS)
Office of Naval Research grant No. N0014-23-1-2051 (HS)
Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University (HS)
Author contributions:
Conceptualization: JS, GK, HS
Methodology: JS, GK, HS
Investigation: JS, GK, HS
Visualization: JS
Supervision: GK, HS
Writing—original draft: JS
Writing—review & editing: JS, GK, HS
Competing interests:
There are no competing interests to declare.
Data and materials availability:
The code and dataset used in the study are publically available on Github: https://github.com/ShaneShang/SimplifiedRPM-Visual-Reasoning.
Supplementary materials
Materials and Methods
Supplementary Text
Figs. S1 to S8
Tables S1 to S2
Supplementary Materials for

### v2

@@ Results
We propose that the trade-off between the generalization signal and D arises from a structured compression effect. All class manifolds—including those for held-out rules—gradually align with a low-dimensional subspace defined by the principal components of the training rules. Across models, the representations of the 35 training rules consistently occupy a low-dimensional space. Notably, we observe a drop in variance explained by the top principal components (PCs) at exactly 353535, matching the number of training rules (Supplementary Fig.4). We hypothesize that the network amplifies the generalization signal by aligning all representations, including those of held-out classes, within this low-dimensional space.
To test this hypothesis, we measured the fraction of variance explained by the top 353535 principal components (PCs) of the training rules for each held-out rule manifold. Higher values indicate more substantial alignment with the low-dimensional subspace of training rules. We observed that as the network signal increases, this alignment becomes more pronounced across models (Fig. 6e-h).
@@
[…]
To assess the effectiveness of the proposed SNR loss, we trained the top-performing SCL model using this objective and evaluated its generalization across the same 151515 rule splits previously used for training with cross-entropy loss. The model effectively optimized the training rule representations, achieving an average SNR of 9.66±0.229.66$\pm$ 0.229.66 ± 0.22 (mean±$\pm$±s.e.m.) across training rule pairs, as evaluated on novel test rows of the training rules. More importantly, the model effectively separated held-out rules, yielding an averaged generalization error of 0.162±0.0180.162$\pm$ 0.0180.162 ± 0.018 (mean±$\pm$±s.e.m.) (Fig. 7a). This performance was statistically indistinguishable from models trained with cross-entropy (CE) loss using one-hot rule labels (Wilcoxon signed-rank test: statistic = 48.048.048.0, $p=0.26$, n.s.), confirming that SNR loss achieves competitive generalization while structuring learned representations according to theoretical predictions.
To further evaluate the efficacy of SNR loss, we compared its performance against Prototypical loss, a widely used loss function designed to learn structured representations that generalize to unseen classes [33]. Prototypical loss operates by embedding inputs into a high-dimensional representation space and forming class prototypes, defined as the centroids of the representations corresponding to each class. The loss function then encourages representations of samples from the same class to cluster around their respective prototypes. Additional details on Prototypical loss are provided in the Supplementary Text. We trained the SCL model using Prototypical loss on the same 151515 rule splits. The resulting model exhibited a generalization error of 0.227±0.0100.227$\pm$ 0.0100.227 ± 0.010, significantly higher than that obtained with SNR loss (Fig. 7a). These findings suggest that while both loss functions promote structured representations, SNR loss is more effective in shaping a geometric representation that enhances generalization.
@@ Discussion
A longstanding challenge in both cognitive science and artificial intelligence is to understand how neural systems generalize abstract relations beyond the rules they were trained on. Here, we take a first step toward addressing this question by introducing a principled geometric framework that links generalization performance to the structure of rule representation. We hypothesize that networks trained to perform reasoning tasks develop structured internal representations even for unseen rules, and these representations form distinct rule manifolds. Building on geometric theory originally developed for object recognition [27], we propose that generalization performance in visual reasoning can be explained by key geometric properties: manifold radii, dimensionality, the distance between manifold centroids, and the overlap between these distances and their subspaces of variation.
To systematically investigate this hypothesis, we introduced SimplifiedRPM, a curated dataset of Raven’s Progressive Matrices tasks that enables systematic evaluation of generalization in both humans and neural networks. We tested a range of models (ResNet [28], WReN [6], ViT [29], and SCL [9]), as well as human participants on the same dataset. Among these, SCL achieved the strongest generalization and most closely mirrored human difficulty rankings across abstract relations. The current work expands on the initial probing of the SCL architecture by providing a quantitative account of compositional generalization and by identifying specific geometric properties that predict generalization performance.
Our analysis revealed that representation geometry accurately predicts generalization ability, and that different models strike different balances between signal and dimensionality. Critically, we observed that not all relational rules are equally generalizable. The best-performing model, SCL, struggled with logic-based rules, which humans also found challenging. Other models, despite high test accuracy on arithmetic rules, failed to generalize them, highlighting a disconnect between testing evaluation and the deeper relational understanding necessary for generalization. Our representation analysis revealed that signal, more than dimensionality or signal-noise overlap, governs these differences across rules. Building on these insights, we propose SNRloss, a theoretically grounded objective function that explicitly optimizes representation geometry. This novel loss function provides adequate training and generalization capabilities.
Layer-wise analyses revealed distinct trajectories in the emergence of relational representations across architectures. In SCL, relational structure emerges abruptly at the first multilayer perceptron (MLP) layer following the convolutional blocks, where global relational information likely becomes accessible through integration across multiple panels. However, this increase does not persist, and the SNR decreased slightly in the subsequent MLP layer. In contrast, Vision Transformers (ViTs) exhibit a more gradual and continuous refinement of relational structure across layers. An intriguing question is whether high-level reasoning operates through discrete, stage-like computations as in SCL, or emerges gradually as in ViTs.
@@
Our findings also refine theoretical predictions about dimensionality. While higher dimensionality can support richer representations and better generalization [27], prior works on object classification also suggest advantages for low-dimensional manifolds in enabling the classification of a greater number of objects [34, 35]. This contrast implies that dimensionality should not be universally maximized or minimized but instead tuned to task demands. Our findings extend this idea by showing that generalization depends not only on dimensionality but mostly on its interaction with pairwise manifold signals. Specifically, we demonstrate that deep networks systematically compress representations into structured, low-dimensional spaces—even for unseen rules. Moreover, this phenomenon is consistently observed across diverse architectures (e.g., ResNet, SCL), indicating that it is a fundamental learning property rather than an artifact of any specific architecture. These findings open new avenues to explicitly leverage this signal versus dimensionality trade-off to improve generalization.
@@
While prior work with Bayesian models accounts for human difficulty patterns across abstract relations by explicitly incorporating prior probabilities for each rule type [36], our findings demonstrate that the SCL model—trained end-to-end without any rule-type-specific assumptions—nonetheless captures human-like patterns of relational difficulty. Furthermore, SCL’s generalization errors closely correlate with human response times. While these correlations certainly do not imply that SCL reasons like humans, our geometric analysis provides a principled foundation for future investigations into the cognitive factors underlying this correlation. The SCL network parallels the brain’s hierarchy: its CNN-based feature extractor mimics the ventral visual cortex, capturing basic features like shape [37, 38, 39, 40], while its later reasoning module reflects higher cortical areas supporting relational reasoning [41, 42, 43, 44]. We propose that the brain encodes rules of input rows through distinct, rule-specific neural manifolds. Our observation of layer-wise changes in the geometry of these rule manifolds offers hypotheses into how neural representations evolve across brain regions [45]. According to the SCL model, early visual areas exhibit high representational dimensionality (D) but low signal, whereas later regions show larger signals with reduced dimensionality. These predictions can be empirically tested by measuring neural activity during abstract reasoning tasks [46, 47, 48].
@@
This study aimed to study the generalization of abstract relations. While the humans in our experiment were not explicitly trained on the rules, their reasoning abilities could be considered "pre-trained" through life experiences, as they may have previously encountered similar rules. In contrast, neural networks were trained under strictly controlled rule splits, ensuring they had no prior exposure to the rules evaluated for generalization. To address this discrepancy, a crucial next step is to develop a new benchmark for human reasoning that incorporates abstract relations unlikely to have been encountered. Such a benchmark would disentangle relational generalization from prior experience, providing a clearer view of their distinct roles in human abstract reasoning. This approach enables a direct comparison between humans and neural networks, offering deeper insights into their generalization mechanisms.
@@
Methods
SimplifiedRPM Dataset
The Raven’s Progressive Matrices (RPM) task [1] is a widely used assessment of abstract relational reasoning, commonly applied in fields such as cognitive psychology and AI evaluation. Each RPM trial presents a 3×3 grid with the bottom-right panel missing. The subject must select the correct answer from eight options to complete the missing panel, ensuring that all three rows or columns in the matrix follow a consistent underlying relational rule.
Inspired by the RPM task, the SimplifiedRPM dataset was specifically designed to assess a model’s ability to generalize abstract relations across different object attributes (Fig. 1). Each row in the dataset follows a unique relational rule that determines how an object’s attribute values change across three panels (Supplementary Fig.5a). The dataset consists of 404040 relational rules, each derived by applying an abstract relation to a specific object attribute. These abstract relations fall into three categories: progression (constant, progression ±2$\pm$ 2± 2, progression ±1$\pm$ 1± 1), arithmetic (addition, subtraction), and logical (AND, OR, XOR). A detailed description of the abstract relations is provided in Supplementary Table 2.
[…]
Dataset generation followed a structured pipeline to ensure consistency and reproducibility. Each row was constructed to enforce a unique rule from our predefined set, eliminating ambiguity. To achieve this, we first sampled attribute values from valid options that conformed to the designated rule. Next, we randomly assigned values to the remaining attributes and resampled if a sampled value inadvertently conformed to an unintended rule. Once attribute values were finalized, they uniquely defined an input row. The corresponding row input was then generated by placing objects with the specified attributes in their designated positions. To maintain dataset balance, we generated 10,000 training rows, 1,000 validation rows, and 1,000 test rows per rule, yielding a total of 48,000 unique rows across all rules (40 rules $\times$ 1,200 samples). Example rows in the dataset are provided in Supplementary Fig.5b-d.
Neural Network Training
To evaluate a model’s ability to generalize abstract relations to unseen object attributes, we implement a rule-splitting strategy in which a subset of rules is withheld during training. Specifically, we randomly select five rules to be held out, ensuring that each abstract relation appears at most once in this set. The remaining 353535 rules are used for model training. After training, the model is evaluated based on its ability to infer the held-out rules. Supplementary Fig.5a provides an example of this rule split, with held-out rules highlighted in yellow and training rules in blue.
During training, models are optimized using supervised learning with a cross-entropy loss function. The models are trained on 10,000 examples per rule across 353535 training rules. Test error (Fig. 1c) is assessed based on classification accuracy using 1,0001,0001 , 000 test examples per rule. These test examples differ from the 10,00010,00010 , 000 training examples used during model training.
To evaluate model performance, we compute the Euclidean distance between the sample row and each of the two choice rows. A trial is considered correct if the correct choice row is closer to the sample row than the incorrect one. The model’s generalization error is then quantified as the proportion of incorrect trials averaged across all 500500500 trials for each of the 202020 rule pairs (Fig. 1e).
We trained four architectures: ResNet-50, WReN, ViT, and SCL. ResNet-50 [28] is a convolutional neural network with residual connections. We employed the standard ResNet-50 architecture without modification. WReN [6] explicitly models pairwise relations between objects and has been previously used for RPM tasks. We optimized the hyperparameters of its relation module and fully connected layers. ViT [29] is a transformer-based model that operates on image patches as input tokens and captures global dependencies via self-attention. We split each of the three panels in a row into a $\times 3$ 3 grid, resulting in 272727 patches per row. We optimized key hyperparameters, including the number of layers, attention heads, and hidden dimensions. Finally, SCL [9] was specifically designed for the RPM task, and we used the original architecture without modifications (Supplementary Fig.6). Detailed model configurations, including layer-wise specifications and hyperparameter tuning, are provided in Supplementary Text.
Training was conducted using stochastic gradient descent with a learning rate of 0.10.10.1, momentum of 0.90.90.9, and weight decay of $5\times 10^{-4}$. To facilitate stable convergence, we applied a learning rate warmup for the first two epochs, followed by a plateau-based scheduler that reduced the learning rate by 0.20.20.2 whenever the validation loss plateaued. Training continued until the classification errors across 1,0001,0001 , 000 validation rows for each of the 353535 training rules stabilized. The final training durations for each model were: ResNet-50 (100100100 epochs), WReN (606060 epochs), ViT (200200200 epochs), and SCL (150150150 epochs). All models were trained using a batch size of 512512512.
Human Experiment
[…]
Before the main experiment, participants read written instructions describing the task and completed three practice trials that introduced relational rules different from the 131313 rule pairs used in the main experiment. Participants received feedback and explanations of the underlying rules during these practice trials. In the main experiment, each participant completed 60 trials randomly selected from a larger pool of 130130130 trials (101010 trials per rule pair), encompassing 131313 distinct rule pairs of varying relational complexity (Supplementary Table 1). Participants were shown images like those in Fig. 1a. The images remained on the screen until the participants made a choice. Participants responded by clicking on one of two choice rows and were allowed unlimited time to provide their answers. No performance feedback was given during the main experiment. Additional information, including illustrative example trials, can be found in Supplementary Fig.7.
Supplementary Information